# Optimization procedure of the baffle of the GroundBIRD Telescope to mitigate stray light

### v1

@@ Abstract
We presented the optimization procedures of the baffle mounted on the GroundBIRD telescope for measuring the polarization of the Cosmic Microwave Background (CMB). The telescope employs dual mirror reflective telescopes installed in a cryostat. The primary objectives were to minimize stray light contamination, maintain the integrity of the main beam, and ensure that thermal loading from the baffle remains significantly below that from the atmosphere. Using quasi-optical simulations, we have optimized the baffle’s aperture angle to suppress stray light without degrading the main beam quality. We confirmed through Moon observations that the optimized baffle design works to eliminate the contamination of the stray light as expected. Furthermore, no measurable degradation in the noise equivalent temperature (NET) was detected, indicating minimal thermal impact. These results show that our baffle optimization strategy effectively reduces systematic errors while maintaining observational sensitivity, providing valuable insights for future CMB experiments with similar optical architectures.
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I Introduction
The precision measurements of the Cosmic Microwave Background (CMB) observations have been playing a central role in exploring the space-time geometry, evolution, and origin of our universe[1, 2]. The current challenge of the CMB observations is high-precision measurements of its polarization signals. One of the key issues in CMB experiments is reducing contamination of systematic errors in the data to achieve the required high accuracy of the measurements. Suppressing the penetration of stray light into the detectors is one of the key subjects for reducing systematics. To prevent contamination of light reflected by the cryostat wall into the detectors, all walls of the cryostat are covered with high-performance millimeter wave absorbers[3, 4, 5]. A ground shield surrounds all the ground-based CMB telescopes, such as BICEP[6] and the Simons Observatory Small Aperture Telescopes[7]. Fig. 1 shows the installation of the GroundBIRD telescope. In some CMB experiments, all the optical components, including telescopes, are enclosed by a box or a cryostat, and a baffle installed in front of the entrance window plays a complementary role for reducing the contamination of the ground signal[6, 7, 8, 9].
@@
In the case of the telescopes adopting reflective mirrors[8, 9, 10, 11, 12, 13], to avoid the light beam disturbance due to the obstacles in the light path, an off-axis reflecting mirror system is employed. To compensate for the instrumental polarization imprinted by the off-axis reflection, the secondary mirror is set appropriately [14]. As shown by [13], two stray light components, which may degrade the accuracy, can reach the detectors in the dual reflecting mirror system. One is reaching the detector without experiencing any reflection by primary and secondary mirrors. It is referred to as the direct component of the stray light. The other is reflected once more from the primary mirror after nominal reflections before reaching the detectors. It is referred to as the spillover component of the stray light. In most of the CMB telescopes employing the reflective mirror system, the primary and secondary mirrors are exposed to the air. In these cases, the focal length of the secondary mirror is stretched by using lenses enclosed by the lens barrel. The lens barrel serves as an absorber, blocking these stray light components. However, in some cases[13, 15, 16, 17] it is difficult to extend the focal length of the secondary mirror because of the limitation of space. In these cases, the baffle plays a role in preventing the penetration of these stray light components into the detectors. To suppress the contamination of the light reflected by the wall of the baffle, the wall of the baffle is also covered by the millimeter wave absorbers.
@@
There are two ways to reduce the contamination of the stray light by optimizing the design of the baffle. One is extending the length. The other is narrowing the aperture diameter of the baffle. In the case of the GroundBIRD experiment, the length of the baffle is restricted by the size of the dome (see Fig. 1). Therefore, narrowing the aperture diameter of the baffle is the only method for reducing the contamination of the stray light. However, narrowing the aperture diameter of the baffle accompanies the risk that the baffle aperture interferes with the main beam. In this study, we present the optimization procedures of the aperture diameter of the baffle for reducing the contamination of the stray light while maintaining the quality of the main beam. We show how the quality of the main beams and stray light rejection depend on the aperture diameter of the baffle of the GroundBIRD based on the results of the simulations and the Moon observations. Although performing the wave optics simulations is ideal for the optimization, we adopt the quasi-optical simulation instead because the physical optics simulation is computationally expensive.
@@
The structure of the paper is as follows. In Section 2, we first introduce the instrumental details of GroundBIRD and its observation strategy. We also introduce the simulation setup. In Section 3, the methodology for optimizing the baffle design is presented. Section 4 presents the details of the optimized baffle design. Section 5 presents the results of the performance verification measurements. Section 6 is devoted to discussion and conclusion.
@@
Figure 1: Photograph of the GroundBIRD telescope installation captured by monitoring camera. The arrangement of ground shield, dome, baffle, and cryostat are shown.
@@ II Methods
II-A The GroundBIRD telescope
The GroundBIRD telescope is a ground-based CMB polarization experiment designed to capture the reionization bump signals imprinted in the large angular scale of approximately 20 degrees in the northern hemisphere for the first time after the Planck satellite[18, 19]. It is considered that it is challenging to reach the reionization bump from ground-based observations due to atmospheric fluctuations. GroundBIRD is attempting to reach the reionization bump through ground-based observations by mitigating atmospheric fluctuations with its unique scan strategy[20]. It performs a rapid rotation scan of up to 20 rotations per minute (RPM) while inclining the telescope 20 to 30 degrees from the zenith. By scanning the large patches of the sky before the sky condition changes significantly, it has the potential to capture the reionization bump.
To achieve the diffraction-limited spatial resolution simultaneously, cutting-edge superconducting detectors, Microwave Kinetic Inductance Detectors (MKIDs), are utilized as the focal plane detectors due to their fast time response. The observed frequency bands are 145 GHz and 220 GHz. Assembly of the detectors is shown in Fig. 2. The central chip mounts 23 pixels for 220 GHz band and the other 6 chips mount 138 pixels for 145 GHz band. GroundBIRD employs the Mizuguchi-Dragone dual reflectors[14]. The whole telescope system is installed in the cryostat and cooled down to 4 K.
The telescope is set at the Teide observatory in Tenerife, Canary Islands. The first light was in 2019[21], and the full detector set was installed in 2023[22].
Figure 2: Photograph of detector array mounted on the focal plane. The pixels for which we show the results of our analysis, are assigned numbers. Pixels numbered from 1 through 8 are selected because their thermal and optical characteristics are most affected by baffle design.
@@ II-B Quasi-optical Simulations with GRASP
We quantified near- and far-field beam patterns for each pixel using quasi-optical simulations. Although we ultimately want to determine how each detector pixel observes the sky—that is, the beam pattern for received light—we instead calculate the beam pattern emitted from each detector pixel. This is justified by the time-reversal invariance of Maxwell’s equations, which ensures that both patterns are identical.
@@
The antenna attached to each MKID first emits the beam. Then, it is modified by the lenslet mounted on each MKID. Up to this stage, the beam pattern is calculated by using a physical optics simulation with CST[23]. Beam patterns after the lenslet are quantified by quasi-optical simulations with GRASP[24]. The results of CST are used for boundary conditions for the GRASP simulations. The setup of the GRASP simulations is shown in Fig. 3a. The apertures of the baffle, the entrance window of the cryostat, and the cold stop are set. The detector pixel position is set according to the design. To set the positions of the primary and the secondary mirrors, the CAD design is incorporated. The mirrors were modeled as perfect conductors. The aperture of the cold stop, which defines the entrance pupil, has a circular shape with a diameter of 220 $\mathrm{mm}$. Parameters describing the characteristics of the baffle are shown in Fig. 3b. To quantify the aperture diameter of the baffle, the opening angle $2\Theta$ is used. The aperture angle, $\Theta$, is defined by the arctangent of the aperture radius divided by the length of the baffle. The baffle length is the distance between the top of the entrance window of the cryostat and the aperture of the baffle. This is the physical length of the baffle plus 10 mm. The angular distance relative to the optical axis of the dual mirror system is denoted by $\theta$.
@@
Figure 3: (a) The simulation setup for the quasi-optical simulations with GRASP. The white holes in the gray areas represent the optical components. The top circular area represents the entrance aperture of the baffle. The entrance window from the cryostat and the aperture of the cold stop, which defines the entrance pupil, are shown by another two circular areas. (b) The definition of the parameters that define the principal geometry of the baffle. (c) A photograph of the wide baffle mounted on the GroundBIRD.
@@
Validations of the simulated beam patterns were performed using the Moon observations, as the aperture size of GroundBIRD (220 mm) is too small to capture the beam pattern, including stray light, using a celestial point source. The specific power observed during Moon observations by detector pixel ID $i$ at frequency of $\nu$ is given by
@@
where $\alpha$ and $\beta$ are spherical polar coordinates of unit sphere centered on the Moon where the origin coincides with the Moon center, $(\alpha_{i},\beta_{i})$ is a direction of the beam center of i-th detector relative to the Moon center, the factor $\frac{1}{2}$ accounts for single polarization detection, $A_{e}$ is the effective aperture area, $t_{r}(\nu)$ is the transmittance of the telescope’s filter including the quantum efficiency of the detector, $U_{i}(\alpha_{i},\beta_{i})$ is the beam pattern of detector pixel $i$, and $I_{\nu}(\alpha,\beta)$ is the intensity distribution of the Moon at frequency of $\nu$. We modeled the Moon’s intensity as a blackbody sphere covered by a dielectric material layer [25] and adapted the model described in [26].
@@ III Baffle Design Optimization
III-A The beam quantification with the wide aperture angle baffle
@@
The beam patterns at $\nu=145$ GHz when using the baffle with the wide aperture angle corresponding to $\Theta=27^{\circ}$(see Fig. 3c) are simulated. In this study, we refer to the baffle with $\Theta=27^{\circ}$ as the “wide baffle” for convenience. Fig. 4a shows the simulated far field 2D beam pattern for the detector pixel denoted by ID 9 as shown in Fig. 2. The dashed line encloses the region where the spillover component appears. We can see some localization of the beam in the central region enclosed by the dashed line. On the other hand, the direct component appears sparsely in the region enclosed by the solid line which is much wider than the region enclosed by the dashed line without any prominent structure. The peak intensity of the direct component is about two order of magnitude less than the peak intensity of the spillover component. Since the spillover component appears closer to the pointing center compared to the direct component, it is expected that the direct component is suppressed automatically when the baffle is designed so as to suppress the spillover component. The specific power when the Moon is picked up by the beam center, $P_{\nu,i,\mathrm{main}}$, is calculated by setting $(\alpha_{i},\beta_{i})=(0,0)$ in equation (1). The specific power when the Moon is picked up by the spillover component, $P_{\nu,i,\mathrm{stray}}$, is calculated by setting $(\alpha_{i},\beta_{i})$ in equation (1) so that the specific power of the Moon picked up by the spillover component becomes maximum. This yields $P_{\nu,i,\mathrm{stray}}/P_{\nu,i,\mathrm{moon}}\sim-34$ dB for pixel ID 9.
@@
Fig. 4b shows a Moon-centered map obtained with the detector pixel ID 9 with the wide baffle. The data were acquired during observations on December 2, 2023. The observation began at 02:52 UTC and lasted for one hour, encompassing one complete observation cycle between the two successive calibration measurements of the resonance frequencies of the MKIDs. The average value of the precipitable water vapor (PWV) was around 2 mm. The inclination angle of the telescope from the zenith was 20 degrees. The telescope was rotated at 10 RPM. The Moon image obtained by the main beam is masked to emphasize the ghost image picked up by the stray light. We observed a bright spot in the bottom left of the map. The relative position of the bright spot relative to the Moon center is consistent with the relative position of the peak of the spillover component to the peak of the main beam shown in Fig. 4a. The Moon observations were conducted multiple times, and we analyzed six observation datasets spanning from December 2, 2023, to October 25, 2024. The sky conditions were similar in these observations: the PWV was around 1.7 mm – 2.5 mm. We confirmed that the bright spot appears at the same location in the maps of all observations. Therefore, we concluded that the bright spot observed in Fig. 4b is the ghost image of the Moon picked up by the spillover component of the stray light. Its peak amplitude compared to that of the Moon center is approximately 25-25 dB. The average level of the background noise is approximately 26-26 dB. Therefore, the net peak amplitude of the ghost image after subtracting the background noise is approximately 30-30 dB. The observed intensity of the ghost image is a few times stronger than that from the beam pattern simulation.
@@
Figure 4: (a) Simulated beam pattern for the pixel ID 9. The relative angular distance from the pointing center, $\theta$, are drawn by concentric circles with 10 degrees interval. The azimuth angles, $\phi$, are drawn by radial lines with 45 degrees intervals. (b) Observed map of the Moon obtained by the the pixel ID 9 when the wide baffle is mounted on the GroundBIRD. Vertical and horizontal axes are the off-set angle relative to the Moon center in the elevation, $\Delta\theta_{\mathrm{M}}$, and azimuth, $\Delta\phi_{\mathrm{M}}$, directions respectively. The maps are masked within a 4-degree radius from the Moon center and divided into grids with 0.1-degree intervals. For each scan, baseline noise was subtracted by fitting a linear function to $\pm$1500 data points from $\Delta\phi_{\mathrm{M}}=0$, while masking the central $\pm$500 data points. The intensity is normalized to the maximum intensity of the Moon image, where the detector response is proportional to $\tan(\frac{\psi}{2})$ and $\psi$ is a phase of fed microwaves for the detector readout[27]. (c) The cross section of the simulated near-field beam pattern for the pixel ID 5.
@@
III-B The baffle aperture angle size dependence of the thermal loading from the baffle

### v2

@@ Abstract
We present the optimization procedure of the baffle mounted on the GroundBIRD telescope for measuring the polarization of the Cosmic Microwave Background (CMB). The telescope employs a dual mirror reflective telescope installed in a cryostat. The primary objectives were to minimize stray light contamination, maintain the integrity of the main beam, and ensure that thermal loading from the baffle remains significantly below that from the atmosphere. Using quasi-optical simulations, we have optimized the baffle’s aperture angle to suppress stray light without degrading the main beam quality. We confirmed through Moon observations that the optimized baffle design works to eliminate the contamination of the stray light as expected. Furthermore, no measurable degradation in the noise equivalent temperature (NET) was detected, indicating minimal thermal impact. These results show that our baffle optimization strategy effectively reduces systematic errors while maintaining observational sensitivity, providing valuable insights for future CMB experiments with similar optical architectures.
@@ I Introduction
The precision measurements of the Cosmic Microwave Background (CMB) observations have been playing a central role in exploring the space-time geometry, evolution, and origin of our universe[1, 2]. The current challenge of CMB observations is high-precision measurements of its polarization signals. One of the key issues in CMB experiments is reducing systematic errors in the data to achieve the required high accuracy of the measurements. Suppressing the penetration of stray light into the detectors is one of the key subjects for reducing systematics. To prevent contamination of light reflected by the cryostat wall into the detectors, all walls of the cryostat are covered with high-performance millimeter wave absorbers[3, 4, 5]. In ground-based CMB experiments, either a fixed ground shield[6, 7] or a co-moving ground shield[8, 9, 10, 11] is applied. In some CMB experiments, a baffle installed in front of the entrance window plays a complementary role in reducing the contamination of the ground signal[6, 7, 12]. Fig. 1 shows the installation of the GroundBIRD telescope.
@@
In the case of telescopes adopting reflective mirrors[11, 12, 13, 14, 15, 16], to avoid diffraction from optical components and their mounting structures in the light path, an off-axis reflecting mirror system is employed. To compensate for the instrumental polarization imprinted by the off-axis reflection, the secondary mirror is positioned to satisfy the Mizuguchi-Dragone conditions, which minimize contamination of instrumental polarization[17]. Two mirror configurations are proposed to satisfy the Mizuguchi-Dragone conditions: the crossed configuration and the Gregorian configuration. The crossed configuration has advantages for realizing a wider field of view and less cross-polarization compared to the Gregorian configuration[18]. However, the crossed configuration is more susceptible to scattering than the Gregorian configuration[18]. The Gregorian configuration, such as SPT[11] and ACT[15], has a real intermediary focus between the secondary and primary mirrors, which can be exploited to effectively shield the focal plane from direct illumination, making it less susceptible to far sidelobes[18]. In contrast, the crossed configuration adopted by GroundBIRD[16] presents a more difficult baffling situation because the focal plane is close to the main beams incident on the primary mirror[18], resulting in higher susceptibility to far sidelobes due to scattering from structures near the focal plane. As shown by [16], two stray light components reach the detectors in the GroundBIRD telescope (see Fig. 2). One is reaching the detectors without experiencing any reflection by the primary and secondary mirrors. It is referred to as the direct component of the stray light. The other is reflected again by the primary mirror after reflection by the secondary mirror before reaching the detectors. It is referred to as the spillover component of the stray light. One way to reduce stray light contamination on the detectors is to increase the distance from the secondary mirror to the detectors by using lenses enclosed in the lens barrel[14]. The lens barrel serves as an absorber, blocking stray light. However, in some experiments[16, 19, 20, 21], it is difficult to increase the distance between the secondary mirror and the detectors due to space limitations. In these cases, a baffle plays a role in preventing penetration of these stray light components into the detectors. To suppress the contamination of the light reflected by the wall of the baffle, the wall of the baffle is also covered by millimeter wave absorbers.
@@
There are two ways to reduce contamination from stray light by optimizing the design of the baffle. One is extending the baffle length. The other is narrowing the top aperture diameter of the baffle. In the case of the GroundBIRD experiment, the baffle length was set to the maximum value allowed by the dome size (see Fig. 1). Therefore, narrowing the top aperture diameter of the baffle is the only method for reducing contamination from stray light. However, narrowing the top aperture diameter of the baffle is accompanied by the risk that the main beams are degraded due to diffraction from the baffle aperture. In this study, we present the optimization procedure for the top baffle aperture diameter to reduce stray light contamination while maintaining the quality of the main beams. We show how the quality of the main beams and stray light rejection depend on the top aperture diameter of the baffle of the GroundBIRD based on the results of the simulations and Moon observations. Although performing wave optics simulations is ideal for the optimization, we adopt the quasi-optical simulation instead because the physical optics simulation is computationally expensive.
@@
The structure of the paper is as follows. In Section II, we first introduce the instrumental details of GroundBIRD and its observation strategy. We also introduce the simulation setup. In Section III, the methodology for optimizing the baffle design is presented. Section IV presents the details of the optimized baffle design. Section V presents the results of the performance verification measurements. Section VI is devoted to discussion and conclusion.
@@
Figure 1: A photograph of the GroundBIRD telescope installation captured by a monitoring camera. The arrangement of the ground shield, dome, baffle, and cryostat is shown.
Figure 2: Primary stray light paths in the GroundBIRD optical system, including key optical components, are illustrated. The main beam is drawn by a solid line. Two types of stray light paths are illustrated: the “spillover” component (dashed line) and the “direct” component (dotted line).
@@ II Methods
II-A GroundBIRD Telescope
The GroundBIRD telescope is a ground-based CMB polarization experiment designed to capture the reionization bump signals imprinted in the large angular scale of approximately 20 degrees in the northern hemisphere for the first time from the ground after the Planck satellite[22, 23]. It is considered that it is challenging to reach the reionization bump from ground-based observations due to atmospheric fluctuations, although the CLASS experiment has conducted observations in the southern hemisphere to contribute to large-scale polarization measurements[24]. GroundBIRD is attempting to reach the reionization bump through ground-based observations by mitigating atmospheric fluctuations with its unique scan strategy[25]. It performs a rapid rotation scan of up to 20 rotations per minute (RPM) while inclining the telescope 20 to 30 degrees from the zenith. By scanning large patches of the sky before the sky condition changes significantly, it has the potential to capture the reionization bump.
To achieve diffraction-limited spatial resolution simultaneously, cutting-edge superconducting detectors, Microwave Kinetic Inductance Detectors (MKIDs), are utilized as the focal plane detectors due to their fast time response. The observed frequency bands are 145 GHz and 220 GHz. Assembly of the detectors is shown in Fig. 3. The central chip mounts 23 pixels for 220 GHz band and the other 6 chips mount 138 pixels for 145 GHz band. GroundBIRD employs Mizuguchi-Dragone dual reflectors[17]. The whole telescope system is installed in the cryostat and cooled down to 4 K.
The telescope is set at the Teide observatory in Tenerife, Canary Islands. First light was in 2019[26], and the full detector set was installed in 2023[27].
Figure 3: A photograph of the detector array mounted on the focal plane. The pixels for which we show the results of our analysis are assigned numbers. Pixels numbered from 1 through 8 are selected because their thermal and optical characteristics are most affected by baffle design.
@@ II-B Quasi-Optical Simulations With GRASP
We quantified near-field and far-field beam patterns for each pixel using quasi-optical simulations. Although we ultimately want to determine how each detector pixel observes the sky—that is, the beam pattern for received light—we instead calculate the beam pattern emitted from each detector pixel. This is justified by the time-reversal invariance of Maxwell’s equations, which ensures that both patterns are identical.
@@
The antenna attached to each MKID first emits the beam. Then, it is modified by the lenslet mounted on each MKID. Up to this stage, the beam pattern is calculated by using a physical optics simulation with CST[28]. Beam patterns after the lenslet are quantified by quasi-optical simulations with GRASP[29]. The results of CST are used for boundary conditions for the GRASP simulations. GRASP simulations are performed for both far-field and near-field beam patterns. The setup for far-field calculations is shown in Fig. 4a, where all optical components including the top aperture of the baffle are incorporated. For near-field calculations, the top aperture of the baffle is excluded to evaluate thermal loading from the baffle wall without diffraction effects. The apertures of the baffle, the entrance window of the cryostat, and the cold stop are set. The detector pixel position is set according to the design. To set the positions of the primary and the secondary mirrors, the CAD model is used. The mirrors are modeled as perfect conductors. The aperture of the cold stop, which defines the entrance pupil, has a circular shape with a diameter of 220 $\mathrm{mm}$. The entrance window from the cryostat also has a circular shape with a diameter of 320 $\mathrm{mm}$. The plane of the cold stop is tilted relative to the entrance window plane to avoid standing waves due to the Fabry-Perot effect. Parameters describing the characteristics of the baffle are shown in Fig. 4b. To quantify the top aperture diameter of the baffle, the opening angle $2\Theta$ is used. The aperture angle, $\Theta$, is defined by the arctangent of the aperture radius divided by the length of the baffle. The baffle length is the distance between the top of the entrance window of the cryostat and the top aperture of the baffle. In our optimization procedure, the baffle length was set to the maximum value (82 cm) allowed by the dome, and only the aperture angle $\Theta$ was varied to suppress stray light while maintaining beam quality. This is the physical length of the baffle plus 10 mm. The angular distance relative to the optical axis of the dual mirror system is denoted by $\theta$.
@@
Figure 4: (a) The simulation setup for the quasi-optical simulations with GRASP. The white holes in the gray areas represent the optical components. The top circular area represents the entrance aperture at the top of the baffle. The entrance window from the cryostat and the aperture of the cold stop, which defines the entrance pupil, are shown by another two circular areas. (b) The definition of the parameters that define the principal geometry of the baffle.
@@
Validations of the simulated beam patterns were performed using Moon observations. Although Jupiter and Venus are observable bright point sources, they are too faint to characterize the beam patterns down to the side lobes within a single-day measurement. Thanks to the wide dynamic range of the MKIDs, it is possible to observe the Moon with GroundBIRD. The Moon is sufficiently bright for characterizing the beam pattern down to the side lobes including stray light with one-hour integration time. Since the Moon has to be treated as an extended source by GroundBIRD, the specific power observed during Moon observations by detector pixel ID $i$ at frequency $\nu$ is given by
@@
where $\alpha$ and $\beta$ are spherical polar coordinates of a unit sphere centered on the Moon where the origin coincides with the Moon center, $(\alpha_{i},\beta_{i})$ is a direction of the beam center of the i-th detector relative to the Moon center, the factor $\frac{1}{2}$ accounts for single polarization detection, $A_{e}$ is the effective aperture area, $t_{r}(\nu)$ is the transmittance of the telescope’s filter including the quantum efficiency of the detector, $U_{i}(\alpha-\alpha_{i},\beta-\beta_{i})$ is the beam pattern of detector pixel $i$, and $I_{\nu}(\alpha,\beta)$ is the intensity distribution of the Moon at frequency $\nu$. We modeled the Moon’s intensity as a blackbody sphere covered by a dielectric material layer [30] and adapted the model described in [31].
@@ III Baffle Design Optimization
III-A Beam Quantification With a Wide Aperture Angle Baffle
Figure 5: A photograph of the wide baffle mounted on the GroundBIRD. The baffle length is 84 cm and the aperture angle $\Theta=27^{\circ}$. (a) External view. (b) Internal view showing the radio-absorptive material (ECCOSORB AN-72) covering the inner wall. The absorber is bonded to the aluminum surfaces using Stycast 2850FTJ adhesive.
@@ III-A Beam Quantification With a Wide Aperture Angle Baffle
The beam patterns at $\nu=145$ GHz when using the baffle with the wide aperture angle corresponding to $\Theta=27^{\circ}$ (see Fig. 5a) were simulated. In this study, we refer to the baffle with $\Theta=27^{\circ}$ as the “wide baffle” for convenience. Fig. 6a shows the simulated far field 2D beam pattern for the detector pixel denoted by ID 9 as shown in Fig. 3. The dashed line encloses the region where the spillover component appears. We can see some localization of the beam in the central region enclosed by the dashed line. On the other hand, the direct component appears sparsely in the region enclosed by the solid line which is much wider than the region enclosed by the dashed line without any prominent structure. The peak intensity of the direct component is about two orders of magnitude less than the peak intensity of the spillover component. Since the spillover component appears closer to the pointing center compared to the direct component, it is expected that the direct component is suppressed automatically when the baffle is designed so as to suppress the spillover component. The specific power when the Moon is picked up by the beam center, $P_{\nu,i,\mathrm{main}}$, is calculated by setting $(\alpha_{i},\beta_{i})=(0,0)$ in (1). The specific power when the Moon is picked up by the spillover component, $P_{\nu,i,\mathrm{stray}}$, is calculated by setting $(\alpha_{i},\beta_{i})$ in (1) so that the specific power of the Moon picked up by the spillover component becomes maximum. This yields $P_{\nu,i,\mathrm{stray}}/P_{\nu,i,\mathrm{moon}}\sim-34$ dB for pixel ID 9.
@@
Fig. 6b shows a Moon-centered map obtained with the detector pixel ID 9 with the wide baffle. The data were acquired during observations on December 2, 2023. The observation began at 02:52 UTC and lasted for one hour, encompassing one complete observation cycle between the two successive calibration measurements of the resonance frequencies of the MKIDs. The average value of the precipitable water vapor (PWV) was around 2 mm. The inclination angle of the telescope from the zenith was 20 degrees. The telescope was rotated at 10 RPM. The Moon image obtained by the main beam is masked to emphasize the ghost image picked up by the stray light. We observed a bright spot in the bottom left of the map. The relative position of the bright spot relative to the Moon’s center is consistent with the relative position of the peak of the spillover component to the peak of the main beam shown in Fig. 6a. Moon observations were conducted multiple times, and we analyzed six observation datasets spanning from December 2, 2023, to October 25, 2024. The sky conditions were similar in these observations: the PWV was around 1.7 mm – 2.5 mm. We confirmed that the bright spot appears at the same location in the maps of all observations. Therefore, we concluded that the bright spot observed in Fig. 6b is the ghost image of the Moon picked up by the spillover component of the stray light. Its peak amplitude compared to that of the Moon center is approximately 25-25 dB. The background noise level is approximately 26-26 dB. Therefore, the net peak amplitude of the ghost image after subtracting the background noise is approximately 30-30 dB. The observed intensity of the ghost image is a few times stronger than that from the beam pattern simulation.
@@
Figure 6: (a) Simulated beam pattern for the pixel ID 9 with the wide baffle ($\Theta=27^{\circ}$). The relative angular distance from the pointing center, $\theta$, is drawn by concentric circles with 10 degrees interval. The azimuth angles, $\phi$, are drawn by radial lines with 45 degrees intervals. The region where the spillover component appears is enclosed by the white dashed line. The region where the direct component appears is enclosed by the white solid line. (b) Observed map of the Moon obtained by the pixel ID 9 when the wide baffle is mounted on the GroundBIRD. Vertical and horizontal axes are the offset angle relative to the Moon center in the elevation, $\Delta\theta_{\mathrm{M}}$, and azimuth, $\Delta\phi_{\mathrm{M}}$, directions respectively. The maps are masked within a 4-degree radius from the Moon’s center and divided into grids with 0.1-degree intervals. For each scan, baseline noise was subtracted by fitting a linear function to $\pm$1500 data points from $\Delta\phi_{\mathrm{M}}=0$, while masking the central $\pm$500 data points. The intensity is normalized to the maximum intensity of the Moon image, where the detector response is proportional to $\tan(\frac{\psi}{2})$ and $\psi$ is a phase of fed microwaves for the detector readout[32].
@@
III-B Baffle Aperture Angle Size Dependence Of the Thermal Loading From the Baffle
Figure 7: The cross section of the simulated near-field beam pattern for the pixel ID 5.